# GSEdit: Efficient Text-Guided Editing of 3D Objects via Gaussian Splatting

### v1

@@ 4 Results
Input shape Prompt IN2N [11] GSEdit (ours)
Refer to caption Refer to caption Refer to $\sim$23 min Refer to $\sim$4 min
Refer to caption Refer to caption Refer to $\sim$20 min Refer to $\sim$4 min
Refer to caption Refer to caption Refer to $\sim$13 min Refer to $\sim$4 min
Figure 6: Qualitative comparison of results for object editing. The outputs of Instruct-NeRF2NeRF by Haque et al. [11] show its inability to focus detail on a single object, hinting at the method only being suitable for entire scenes. We provide additional results in comparison with other methods in the supplementary materials.
4.1 Qualitative Results
Our qualitative results are shown in Figures 1, 3 and 6. GSEdit shows remarkable capabilities in consistently editing Gaussian splatting object so that the final result can be extracted as a mesh. It is possible to modify the shape of objects as well as the color and yield good results once the correct parameters are set. However, the editing power is heavily bound to the capabilities of IP2P (as we discuss later in 5), so finding good parameters is fundamental to achieving the desired mesh as output. GSEdit is able to modify the shape, the color, and the overall style of the input meshes according to the input prompt. Our experiments include changes in object identity with preservation of original features, and in visual appearance, such as material or artistic style.
4.2 Quantitative Results
Text-guided editing is mostly a subjective task, in the sense that two distinct “evaluators” could expect different outcomes out of the same prompt: even the idea of a benchmark set is ill-defined in this setting. As a consequence, it is nontrivial to identify some metric that objectively expresses how accurate the modifications for given inputs are. Following previous work [11, 27, 19], we analyze the edited object in the CLIP [28] space to compare it with the prompt. In particular, the two following CLIP-based metrics can be used.
@@ 4.2 Quantitative Results
This metric compares a render of the original 3D object $x^{(i)}$ from some viewpoint $i$ the and the render of the edited object $\hat{x}^{(i)}$ from the same viewpoint and two prompts, one describing the original object $T^{(i)}$ (e.g. “A cartoonish penguin”) and describing the edited object $\hat{T}^{(i)}$(e.g. “A cartoonish penguin made of gold”). Then, we embed the image pair and the text pair in the CLIP space, and we compute the cosine similarity between the offsets $\left(x_{\text{CLIP}}^{(i)}-\hat{x}_{\text{CLIP}}^{(i)}\right)$ and $\left(T_{\text{CLIP}}^{(i)}-\hat{T}_{\text{CLIP}}\right)$, as:
$\text{CLIP}_{sim}=\dfrac{\left(x_{\text{CLIP}}^{(i)}-\hat{x}_{\text{CLIP}}^{(i)}\right)\cdot\left({T_{\text{CLIP}}^{(i)}-\hat{T}_{\text{CLIP}}^{(i)}}\right)}{\lVert x_{\text{CLIP}}^{(i)}-\hat{x}_{\text{CLIP}}^{(i)}\rVert\cdot\|T_{\text{CLIP}}^{(i)}-\hat{T}_{\text{CLIP}}^{(i)}\|}$ (11)
@@
In addition to evaluating the directional similarity, assessing the consistency of the edited object across camera views is imperative. To quantify this, we focus on analyzing pairs of adjacent frames, capturing the original and edited 3D objects from consecutive viewpoints along a camera path, denoted by $x^{(i)},x^{(i+1)}$ for the original, and $\hat{x}^{(i)},\hat{x}^{(i+1)}$ for their edited counterparts. Embedding these frames within the CLIP space allows us to easily evaluate the alignment of the editing across them. As for directional similarity, we obtain the CLIP directional consistency metric proposed in [11] as the cosine similarity between the change in the CLIP embeddings from one frame to the next:
@@
$\text{CLIP}_{cons}=\dfrac{\left(\hat{x}_{\text{CLIP}}^{(i)}-x_{\text{CLIP}}^{(i)}\right)\cdot\left({\hat{x}_{\text{CLIP}}^{(i+1)}-x_{\text{CLIP}}^{(i+1)}}\right)}{\lVert\hat{x}_{\text{CLIP}}^{(i)}-x_{\text{CLIP}}^{(i)}\rVert\cdot\|{\hat{x}_{\text{CLIP}}^{(i+1)}-x_{\text{CLIP}}^{(i+1)}}\|}$ (12)
Results for both metrics and execution time are reported in Table 1. Additional results are provided in the supplementary material.
Table 1: Evaluation of quality and performance metrics for our method and Instruct-NeRF2NeRF. While the two methods are comparable for view-wise consistency, GSEdit shows notable benefits with respect to the coherence of the edited scene with the given prompts. However, our method is extremely more efficient, only requiring 16% of the time to run from start to finish on average.
GSEdit (ours) IN2N [11]
$\text{CLIP}_{sim}(\uparrow)$ 0.2345 0.1459
$\text{CLIP}_{cons}(\uparrow)$ 0.8147 0.8067
Time (s) 239 1493
Refer to caption Refer to caption Refer to caption
@@
Figure 7: Illustration of Instruct-Pix2Pix limitations: Observed from two distinct angles, the tiger consistently appears to face the camera due to IP2P’s processing. Such uniform orientation across various viewpoints leads to a blurred effect in the eye region upon reconstruction.
5 Conclusions
@@
We introduced GSEdit, a method for efficient text-guided editing of 3D objects based on image diffusion models and the Gaussian splatting scene representation. Our method is simple, building on top of previous research, and it allows to solve a complex problem much more efficiently than previous contributions, without sacrificing the quality of its output: in fact, our model yields qualitatively superior outputs to state of the art methods for text-guided editing of entire scenes (rather than focused on 3D objects), but it does so in a fraction of the time. As we previously outlined, our contribution has significant applications in the fields of 3D vision and graphics, and it leaves room for improvement in future research: one particularly interesting avenue would be to find new ways to increase the view-wise consistency of the diffusion model’s edits of the current scene views.

### v2

@@ 4. Results
4.1. Qualitative Results
Our qualitative results are shown in Figures 1, 3, 7 and 6. GSEdit shows remarkable capabilities in consistently editing Gaussian splatting object so that the final result can be extracted as a mesh. It is possible to modify the shape of objects as well as the color and yield good results once the correct parameters are set. However, the editing power is heavily bound to the capabilities of IP2P (as we discuss later in Section 5), so finding good parameters is fundamental to achieving the desired mesh as output. GSEdit is able to modify the shape, the color, and the overall style of the input meshes according to the input prompt. Our experiments include changes in object identity with preservation of original features, and in visual appearance, such as material or artistic style.
4.2. Quantitative Results
Text-guided editing is mostly a subjective task, in the sense that two distinct “evaluators” could expect different outcomes out of the same prompt: even the idea of a benchmark set is ill-defined in this setting. As a consequence, it is nontrivial to identify some metric that objectively expresses how accurate the modifications for given inputs are. Following previous work (Haque et al., 2023; Qian et al., 2024; Liu et al., 2023), we analyze the edited object in the CLIP (Radford et al., 2021) space to compare it with the prompt. In particular, the two following CLIP-based metrics can be used.
@@ 4.2. Quantitative Results
This metric compares a render of the original 3D object $x^{(i)}$ from some viewpoint $i$ the and the render of the edited object $\hat{x}^{(i)}$ from the same viewpoint and two prompts, one describing the original object $T^{(i)}$ (e.g. “A cartoonish penguin”) and describing the edited object $\hat{T}^{(i)}$(e.g. “A cartoonish penguin made of gold”). Then, we embed the image pair and the text pair in the CLIP space, and we compute the cosine similarity between the offsets $\left(x_{\text{CLIP}}^{(i)}-\hat{x}_{\text{CLIP}}^{(i)}\right)$ and $\left(T_{\text{CLIP}}^{(i)}-\hat{T}_{\text{CLIP}}^{(i)}\right)$, as:
(11) $\text{CLIP}_{\textit{sim}}=\dfrac{\left(x_{\text{CLIP}}^{(i)}-\hat{x}_{\text{CLIP}}^{(i)}\right)\cdot\left({T_{\text{CLIP}}^{(i)}-\hat{T}_{\text{CLIP}}^{(i)}}\right)}{\left\lVert x_{\text{CLIP}}^{(i)}-\hat{x}_{\text{CLIP}}^{(i)}\right\rVert\cdot\left\lVert T_{\text{CLIP}}^{(i)}-\hat{T}_{\text{CLIP}}^{(i)}\right\rVert}$
@@
In addition to evaluating the directional similarity, assessing the consistency of the edited object across camera views is imperative. To quantify this, we focus on analyzing pairs of adjacent frames, capturing the original and edited 3D objects from consecutive viewpoints along a camera path, denoted by $x^{(i)},x^{(i+1)}$ for the original, and $\hat{x}^{(i)},\hat{x}^{(i+1)}$ for their edited counterparts. Embedding these frames within the CLIP space allows us to easily evaluate the alignment of the editing across them. As for directional similarity, we obtain the CLIP directional consistency metric proposed by Haque et al. (2023) as the cosine similarity between the change in the CLIP embeddings from one frame to the next:
@@
(12) $\text{CLIP}_{\textit{cons}}=\dfrac{\left(\hat{x}_{\text{CLIP}}^{(i)}-x_{\text{CLIP}}^{(i)}\right)\cdot\left({\hat{x}_{\text{CLIP}}^{(i+1)}-x_{\text{CLIP}}^{(i+1)}}\right)}{\left\lVert\hat{x}_{\text{CLIP}}^{(i)}-x_{\text{CLIP}}^{(i)}\right\rVert\cdot\left\lVert{\hat{x}_{\text{CLIP}}^{(i+1)}-x_{\text{CLIP}}^{(i+1)}}\right\rVert}$
Additionally, we report $\text{CLIP}_{\textit{text}}$, the CLIP text similarity, which evaluates to the cosine similarity between the CLIP embeddings of the generative prompt (i.e., “a mammoth” rather than “turn it into a mammoth”) and the CLIP embeddings of the edited object render. Results for all metrics and execution time are reported in Table 1.
Table 1. Evaluation of quality and performance metrics for our method and multiple baselines: IN2N and GaussianEditor are scene editing methods, while Vox-E is focused on 3D objects, like our method. GSEdit presents a clear superiority in terms of the $\text{CLIP}_{\textit{sim}}$ metric (2$\sim$ 2$\times$ 2 × with respect to all baselines) and runtime (speedups of 6.25×,1.65×,13.40×6.25$\times$,1.65$\times$,13.40$\times$6.25 × , 1.65 × , 13.40 × respectively). This highlights the efficiency and the coherence of the editing provided by our method. For metrics where it does not exceed the baselines scores, it still provides competitive results, while the same does not hold for the other three methods.
IN2N GaussianEditor Vox-E GSEdit
$\text{CLIP}_{\textit{sim}}(\uparrow)$ 0.1382 0.1313 0.1196 0.2345
$\text{CLIP}_{\textit{cons}}(\uparrow)$ 0.8226 0.8436 0.6500 0.8147
$\text{CLIP}_{\textit{text}}(\uparrow)$ 0.1986 0.2236 0.2868 0.2631
Time (s) 1493 394 3204 239
@@
Figure 8. Illustration of Instruct-Pix2Pix limitations: Observed from two distinct angles, the tiger consistently appears to face the camera due to IP2P’s processing. Such uniform orientation across various viewpoints leads to a blurred effect in the eye region upon reconstruction.
5. Conclusions
@@
We introduced GSEdit, a method for efficient text-guided editing of 3D objects based on image diffusion models and the Gaussian splatting scene representation. Our method is simple, building on top of previous research, and it allows to solve a complex problem much more efficiently than previous contributions, without sacrificing the quality of its output: in fact, our model yields qualitatively superior outputs to state of the art methods which, rather than focusing on 3D objects, tackle the problem for entire scenes, and it does so in a fraction of the time. While a baseline specifically for 3D object editing provides more competitive results to our method, the difference in editing time in this case increases dramatically, proving the usefulness of our contribution. As we previously outlined, our contribution has significant applications in the fields of 3D vision and graphics, and it leaves room for improvement in future research: one particularly interesting avenue would be to find new ways to increase the view-wise consistency of the diffusion model’s edits of the current scene views.